North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Home Locomotives Q6 63395 (NER T2 2238) Q6 Blog Saturday 21st October 2017

Saturday 21st October 2017

A good turn out yesterday with Trevor Wilford, Peter Ellis, Les Harper, Nigel Hall, Bill Dobson and Ian Pearson. Nigel brought in a new supply of biscuits and Peter shared some of his mince pies with us at lunch time - the treats you are missing by not being there! Not even any mice at the moment to share them with.

After the usual cuppa, work started with Les and Trevor set about trying to remove the sheared stud from the cab floor which is part of the fall plate fixing on the R/H side. This task took all day by drilling and tapping with a chisel: we even borrowed some easy out taps from the MPD but they did not work. Les persevered and eventually the shell of metal came out and we could get an (almost) plug tap down the blind hole. The thread was quite badly worn by this time and the ability of those threads to hold in even a new bolt are questionable. We may need to resort to locktight to keep that bolt in place and rely upon the other three bolts to hold the hinge down. One lesson is that we could do with a proper tap wrench at Grosmont. Les, Trevor and Nigel then fitted the fall plate: Trevor straightened it. The sharp edge at the back RHS needs dressing on Wednesday.

Ian removed the left leading axle box top plate for drilling lubrication holes but was unable to remove the right leading plate as it is trapped by a piece of packing on top of the axle box. Ian and Peter wound up the spring hanger nuts on the right leading spring so as to take the weight off the packing but this did not work. Looks like a jack will need to be used next week. Trevor has welded nuts above the holes on five of the axle box top plates so corks can now be screwed into the nuts and trimmings made and fitted to the axle boxes.

Bill and Nigel did some measuring of small tube lengths but this was made more difficult as quite a number of tubes at the bottom were blocked with ash, as has previously been reported. A rod was brought up from the MPD and one was cleaned out and a correct length was achieved.

Bill and Ian had to leave early about 14.30, but the rest of the gang continued working till 16.00 and finished off Peter's mince pies.

There continued to be a steady stream of visitors through the day, although with the end of the operating season next weekend they will probably become an endangered species.